OpenStudy (anonymous):

If the m<1 + m<2 + m<3 = 180 and m<3 = 90, what property of equality justifies the conclusion that m<1 + m<2 = 90 ? A. reflexive property B. transitive property C. symmetric property D. subtraction property btw, the < sign is actually an angle.

OpenStudy (cgreenwade2000):

Review the properties here http://www.regentsprep.org/Regents/math/geometry/GPB/theorems.htm

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

B.

OpenStudy (cgreenwade2000):

Transitive says If a = b and b = c, then a = c.

OpenStudy (cgreenwade2000):

What happened from here m1+m2+m3 = 180 m3-90 to here m1+m2=90

OpenStudy (anonymous):

so its ike if m<1 + m<2 + m<3 = 180, and m<3 = 90, then m<1 + m<2 = 90

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

the leap?

OpenStudy (cgreenwade2000):

>_> What mathematical operation happened between m1+m2+m3=180 to m1+m2=90

OpenStudy (anonymous):

subtraction?

OpenStudy (cgreenwade2000):

Something is missing is the second one. How did it go missing? Correct! Subtraction

OpenStudy (cgreenwade2000):

We subtracted m3 which is 90 on both sides.

OpenStudy (cgreenwade2000):

So what property do we know of that involves subtracting things on both sides to keep the equation equal.

OpenStudy (anonymous):

the subtraction property of equality